A company is projected to generate free cash flows of $329 million next year, growing at a 5.7% rate until the end of year 3. After that, cash flows are expected to grow at a stable rate of 2.1%. The company's cost of capital is 13.3%. The company owes $171 million to lenders and has $37 million in cash. If it has 114 million shares outstanding, what is your estimate for its stock price? Round to one decimal place.
Answer #1

Given for a company,

FCFF1 = $329 million

growth rate for next 3 years = 5.7%

So, FCFF2 = 329*1.057 = $347.75 million

FCFF3 = 347.75*1.057 = $367.57 million

FCFF4 = 367.57*1.057 = $388.53 million

After that, cash flows are expected to grow at a stable rate of 2.1%.

So, FCFF5 = 388.53*1.021 = $396.69 million

Cost of capital Kc = 13.3%

So terminal value of firm at year 4 is

V4 = FCFF5/(Kc-g) = 396.69/(0.133-0.021) = $3541.84 million

So value today is sum of PV of future cash flows and PV of value at year 4

So, V0 = 329/1.133 + 347.75/1.133^2 + 367.57/1.133^3 + 388.53/1.133^4 + 3541.84/1.133^4 = $3199.15

So value of firm today = $3199.15 million

MV of debt = $171 million

Cash = $37 million

number of shares outstanding = 114 million

we know that value of firm = MV of debt + Current market price of stock * no. of share outstanding - cash

so, 3199.15 = 171 + CMP*114 - 37

CMP = $26.89

estimate for its stock price is $26.89